# How Transformers Implement Induction Heads: Approximation and Optimization Analysis

### v2

@@ Abstract
Transformers* Equal contribution. have demonstrated exceptional in-context learning capabilities, yet the theoretical understanding of the underlying mechanisms remain limited. A recent work (Elhage et al., 2021) identified a “rich” in-context mechanism known as induction head, contrasting with “lazy” $n$-gram models that overlook long-range dependencies. In this work, we provide both approximation and optimization analyses of how transformers implement induction heads. In the approximation analysis, we formalize both standard and generalized induction head mechanisms, and examine how transformers can efficiently implement them, with an emphasis on the distinct role of each transformer submodule. For the optimization analysis, we study the training dynamics on a synthetic mixed target, composed of a 4-gram and an in-context 2-gram component. This setting enables us to precisely characterize the entire training process and uncover an abrupt transition from lazy (4-gram) to rich (induction head) mechanisms as training progresses.
@@ 1 Introduction
In this paper, we aim to provide a comprehensive analysis of how transformers can efficiently implement induction heads. The first key problem is to rigorously formalize induction heads and evaluate the efficiency of transformers in representing them. According to Elhage et al. (2021), the original induction head can be implemented using a two-layer, twelve-head transformer without feed-forward networks (FFNs). However, practical scenarios demand more powerful induction heads. Thus, it is crucial to generalize the mechanism behind and explore how different transformer submodules, such as varying the number of attention heads or incorporating FFNs, impact the transformer’s ability to implement them. This forms our first research objective:
@@
(Optimization). Understand how transformers transition from relying on $n$-gram patterns to employing the induction head mechanism as training progresses.
@@
  • Optimization analysis: how learning undergoes a sharp transition from $n$-gram to induction head. We study the learning dynamics of a two-layer transformer without FFNs for a mixed target, composed of a 4-gram and an in-context 2-gram component. This toy setting allows us to capture the entire training process precisely. Specifically, we show that learning progresses through four phases: partial learning of the 4-gram, plateau of induction head learning, emergence of the induction head, and final convergence, showcasing a sharp transition from 4-gram to induction head. Our analysis identifies two key drivers of the transition: 1) time-scale separation due to low- and high-order parameter dependencies in self-attention, and 2) speed differences caused by the relative proportions of the two components in the mixed target. Additionally, in our analysis, we introduce a novel Lyapunov function that exploits the unique structure of self-attention, which may be of independent interest.
2 Related Works
Expressiveness of transformers. Theoretically, Dehghani et al. (2019); Pérez et al. (2021); Wei et al. (2022) explored the Turing-completeness of transformers; Yun et al. (2019) established the universal approximation property of transformers. Subsequent studies examined the efficiency of transformers in representing specific functions or tasks, such as sparse functions (Edelman et al., 2022), targets with nonlinear temporal kernels (Jiang and Li, 2023), practical computer programs (Giannou et al., 2023), long but sparse memories (Wang and E, 2024), induction head (Sanford et al., 2024a; b; Rajaraman et al., 2024), and memorization and reasoning (Chen and Zou, 2024). Besides, many studies suggest that transformers achieve in-context learning by approximating gradient-based iterations across various layers (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023; Mahankali et al., 2023; Bai et al., 2023; Shen et al., 2023). Besides, several studies explored the limitation of transformer’s expressivity, particularly in modeling formal languages or simulating circuits (Hahn, 2020; Weiss et al., 2021; Bhattamishra et al., 2020; Merrill et al., 2022; Merrill and Sabharwal, 2023). Among all these works, the most closely related to ours are Rajaraman et al. (2024), which examined a generalized induction head similar to our Eq. (6). Specifically, they showed that multi-layer transformers with single-head attention can implement this mechanism. In contrast, we prove that two-layer transformers are sufficient if multihead attention is used.
Training dynamics of transformers. To gain insights into the dynamics of training transformers, several studies have analyzed simplified transformers on toy tasks. These tasks include learning distinct/common tokens (Tian et al., 2023), leaning balance/inblanced features (Huang et al., 2023), linear regression task (Zhang et al., 2023; Ahn et al., 2024), multi-task linear regression (Chen et al., 2024a), binary classification (Li et al., 2024), Transformer with diagonal weights (Abbe et al., 2024), learning causal structure (Nichani et al., 2024), sparse token selection task (Wang et al., 2024), and learning $n$-gram Markov chain (Chen et al., 2024b). Additionally, studies such as Ataee Tarzanagh et al. (2023); Tarzanagh et al. (2023); Vasudeva et al. (2024) have analyzed settings where transformers converge to margin maximization solutions and Thrampoulidis (2024) examined the implicit bias of next-token prediction. Among these works, the most closely related to ours are Nichani et al. (2024) and Chen et al. (2024b), which proved that two-layer transformers can converge to induction head solutions. In this work, we explore a setting where the target is a mixture of 4-gram and induction head. We show that two-layer transformers can effectively converge to this mixed target and provide a precise description of the learning process associated with each component. Importantly, we are able to capture the abrupt transition from learning 4-gram patterns to mastering the induction head mechanism—a critical phase in the learning of induction heads, as highlighted in previous studies (Elhage et al., 2021; Olsson et al., 2022).
@@ 3 Preliminaries
Notations. For $k\in\mathbb{N}^{+}$, let $[k]=\{1,2,\dots,k\}$. For a vector $v$, we denote by $\left\|v\right\|_{p}$ the $\ell_{p}$ norm, where $1\leq p\leq\infty$. For a matrix $A=(a_{i,j})$, we denote by $\|A\|$, $\|A\|_{F}$ the spectral and Frobenius norms, respectively; let $\|A\|_{1,1}=\sum_{i,j}|a_{i,j}|$. We use standard big-O notations $\mathcal{O},\Omega,\Theta$ to hide absolute positive constants, and use $\tilde{\mathcal{O}},\tilde{\Omega},\tilde{\Theta}$ to further hide logarithmic constants.
Sequence modeling. Given a sequence of tokens $(x_{1},x_{2},x_{3},\dots)$ with each token lying in $\mathbb{R}^{d}$, let $X_{L}=(x_{1},x_{2},\dots,x_{L})\in\mathbb{R}^{d\times L}$ and $X_{m:n}=(x_{m},x_{m+1},\dots,x_{n})\in\mathbb{R}^{d\times(n-m+1)}$. Then, we consider the next-token prediction task: predict $x_{L+1}$ using $X_{L}=(x_{1},x_{2},\dots,x_{L})$.
@@
Transformer is designed to more efficiently capture long-range dependencies in sequence modeling (Vaswani et al., 2017). Specifically, given an $L$-token input sequence $X=(x_{1},\cdots,x_{L})\in\mathbb{R}^{d\times L}$, an $U$-layer transformer TF processes it as follows. First, each input token is embedded into a higher-dimensional space through an embedding layer:
$x_{s}^{(0)}=W_{E}x_{s}+b_{E},\ s\in[L],\ \text{ with }W_{E}\in\mathbb{R}^{D\times d},b_{E}\in\mathbb{R}^{D}.$
Next, the $U$-layer attention blocks process the embedded sequence $X^{(0)}=(x_{1}^{(0)},\cdots,x_{L}^{(0)})$ as follows, and the output of the final layer is taken as the output sequence ${\textnormal{{TF}}}(X)=X^{(L)}\in\mathbb{R}^{D\times L}$:
@@
Here, $\textnormal{{FFN}}^{(u)}$ denotes a (point-wise) two-layer FFN of width $M$, and ${\textnormal{{SA}}}^{(u)}$ represents the multi-head self-attention operation. Specifically, when applied to a sequence $Z=(z_{1},\cdots,z_{L})\in\mathbb{R}^{D\times L}$, ${\textnormal{{SA}}}^{(l)}$ operates it as follows:
${\textnormal{{SA}}}^{(u)}(Z)=$ $W_{O}^{(u)}\sum_{h=1}^{H_{u}}{\textnormal{{SA}}}^{(u,h)}(Z),$ (2)
${\textnormal{{SA}}}^{(u,h)}(Z)=$ $\left(W_{V}^{(u,h)}Z\right){\rm softmax}\left(\left\langle W_{Q}^{(u,h)}Z,W_{K}^{(u,h)}Z\right\rangle+R^{(u,h)}\right),$
@@
Relative positional encoding (RPE). In this work, we focus on the Alibi RPE (Press et al., 2022), where $R_{ij}^{(u,h)}$ follows a Toeplitz structure, i.e. $R_{ij}^{(u,h)}=\phi(i-j;p^{(u,h)})$ for $i,j\in[L]$. Here, $p^{(u,h)}$’s are learnable parameters and we consider $\phi(\cdot;p)$ of the following form:
@@
Note that we adopt Alibi RPE only for simplicity and our results can be easily extended to other additive RPEs, such as T5 (Raffel et al., 2020).
@@ 4.1 Vanilla Induction Heads
Figure 1: An illustrative example of the induction head (taken from Elhage et al. (2021)). A pre-trained two-layer eight-head transformer without FFNs proceeds the context [$\cdots$. The D] by retrieving the preceding information most relevant to the current token (D), then copying and pasting the subsequent token (the green urs) as the current prediction.
@@
This model retrieves previous tokens $x_{s-1}$’s that are similar to the current token $x_{L}$ based on a dot-product similarity, and then copies and pastes $x_{s-1}$’s subsequent token $x_{s}$ as the current prediction $x_{L+1}$. The softmax normalization is applied to the similarity scores to ensure the sparsity of relevant information. For example, in Figure 1, the current token $x_{L}$ is D, and the model retrieves previous tokens similar to D, copying and pasting its subsequent token urs as the prediction.
@@
Let $\textnormal{{IH}}_{2}$ satisfy Eq. (4). Then exists a constant $C>0$ and a two-layer single-head transformer TF (without FFNs), with $D=2d$, $W_{K}^{(1,1)}=W_{Q}^{(1,1)}=0$, $p^{(2,1)}=0$, and $\|W_{K}^{(2,1)}\|,\|W_{Q}^{(2,1)}\|\leq\mathcal{O}(1,\|W^{\star}\|_{F})$, such that
@@
It is worth noting that the approximation efficiency is independent of the sequence length. The proof is provided in Appendix A.1, offering the following insights into how two-layer single-head transformers without FFNs implement vanilla induction heads:
  • In the first layer, RPE in SA is used to approximate the memory kernel: $\rho(\cdot)=\mathbb{I}\{\cdot=1\}$. Thus, SA can capture the preceding token via $x_{s-1}=\sum_{j\geq 1}x_{s-j}\rho(j)$ for each token $x_{s}$, which is crucial for future comparison. Notably, DP in this layer is not essential and can be omitted.
  • In the second layer, DP is used to calculate the similarity $\left\langle\cdot,W^{\star}\cdot\right\rangle$. Then SA can compare the current token $x_{L}$ with all previous tokens to identify those tokens that are similar to $x_{L}$. If a token is found to be similar to $x_{L}$, the model copies its subsequent token and pastes it as the current prediction. Notably, RPE in this layer is not essential and can be omitted.
@@
Our theoretical analysis is consistent with the experimental observations reported in Elhage et al. (2021). Specifically, the experiments there demonstrate that SA in the first layer attends to adjacent tokens, while SA in the second layer retrieves information related to the current token. Our analysis identifies components responsible for these two operations, and reveals that single-head transformers suffice to perform them efficiently.
@@ 4.2 Generalized Induction Heads: In-context $n$-gram and Generic Similarity
$\textnormal{{IH}}_{n}(X_{L})=\sum_{s=n}^{L-1}{\rm softmax}\Big{(}X_{L-n+2:L}^{\top}W^{\star}X_{s-n+1:s-1}\Big{)}x_{s},$ (6)
where the patch $X_{s-n+1:s-1}$ incorporates richer local information near $x_{s}$ and $X_{L-n+2:L}$ denotes the current patches. This formulation is more general than Eq. (4), which only focuses on $x_{s-1}$. This induction head operates based on the similarity between the $n$-gram pairs: $(X_{s-n+1:s-1};X_{L-n+2:L})$ for $s=n,\dots,L-1$.
Integrating richer local information facilitates more accurate information retrieval. The model (6) retrieves previous $(n-1)$-token patch that are similar to the current $(n-1)$-token patch, thereby generalizing the vanilla induction head (4), which considers only single-token retrieval. For example, as depicted in Figure 1, if the current local information is The D (comprising two tokens), and prior local information such as Mr D and Mrs D is identified as similar to The D, transformer would copy and paste their subsequent token, urs, as the prediction.
@@
Let $\textnormal{{IH}}_{n}$ satisfy Eq. (6). Then for any $q\in\mathbb{N}^{+}$ , there exists a constant $C_{q,n}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ (without FFNs), with $D=nd$, such that:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq\frac{C_{q,n}}{H^{q}}.$
This theorem demonstrates that two-layer multi-head transformers, even without FFNs, can efficiently implement the generalized induction head (6). Notably, the approximation error scales as $\mathcal{O}(H^{-q})$, where $q$ can be arbitrarily large. The proof of this theorem is provided in Appendix A.2.
@@
$\textnormal{{GIH}}_{n}(X_{L})=\sum_{s=n}^{L-1}{\rm softmax}\Big{(}g\big{(}X_{L-n+2:L},X_{s-n+1:s-1}\big{)}\Big{)}x_{s},$ (7)
@@
Let $\textnormal{{GIH}}_{n}$ satisfy Eq. (7). Suppose the similarity function $g$ is $\alpha$-well-behaved (see Definition A.7). Then, for any $q\in\mathbb{N}^{+}$, there exist constants $A_{g,q,n},B_{g,\alpha}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ with FFNs of width $M$, such that
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq\frac{A_{g,q,n}}{H^{q}}+\frac{B_{g,\alpha}L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}}.$
@@
The role of FFNs. In contrast to Theorem 4.3, transformer models in Theorem 4.4 include FFNs. These FFN layers are used to approximate the similarity function $g$. Specifically, we consider the proper orthogonal decomposition (POD) of $g$, which can be viewed as an extension of the matrix singular value decomposition (SVD) applied to functions of two variables. For $g:\mathcal{I}\times\mathcal{I}\to\mathbb{R}$, its POD is $g(u,v)=\sum_{k=1}^{\infty}\sigma_{k}\phi_{k}(u)\psi_{k}(v)$, where $\phi_{k},\psi_{k}$ are orthonormal bases for $L^{2}(\mathcal{I})$ (see Appendix D for details). Intuitively, the FFN in the first layer is used to efficiently approximate $K$ bases ($\phi_{i}$’s and $\psi_{i}$’s). Then, in the second layer, DP in SA can approximately reconstruct $g$ by using the truncated sum $g(u,v)\approx\sum_{k=1}^{K}\sigma_{k}\phi_{k}(u)\psi_{k}(v)$. The complete proof is deferred to Appendix A.3.

### v3

# How Transformers Get Rich: Approximation and Dynamics Analysis

@@ Abstract
Transformers have demonstrated exceptional in-context learning capabilities, yet the theoretical understanding of the underlying mechanisms remains limited. A recent work (Elhage et al., 2021) identified a “rich” in-context mechanism known as induction head, contrasting with “lazy” $n$-gram models that overlook long-range dependencies. In this work, we provide both approximation and dynamics analyses of how transformers implement induction heads. In the approximation analysis, we formalize both standard and generalized induction head mechanisms, and examine how transformers can efficiently implement them, with an emphasis on the distinct role of each transformer submodule. For the dynamics analysis, we study the training dynamics on a synthetic mixed target, composed of a 4-gram and an in-context 2-gram component. This controlled setting allows us to precisely characterize the entire training process and uncover an abrupt transition from lazy (4-gram) to rich (induction head) mechanisms as training progresses.
@@ 1 Introduction
In this paper, we aim to provide a theoretical analysis of how transformers can efficiently implement induction heads. The first key problem is to rigorously formalize induction heads and evaluate the efficiency of transformers in representing them. According to Elhage et al. (2021), the original induction head can be implemented using a two-layer, twelve-head transformer without feed-forward networks (FFNs). However, practical scenarios demand more powerful induction heads. Thus, it is crucial to generalize the mechanism behind and explore how different transformer submodules, such as varying the number of attention heads or incorporating FFNs, impact the transformer’s ability to implement them. This forms our first research objective:
@@
(Dynamics). Understand how transformers transition from relying on $n$-gram patterns to employing the induction head mechanism as training progresses.
@@
  • Dynamics analysis: how learning undergoes a sharp transition from $n$-gram to induction head. We study the learning dynamics of a two-layer transformer without FFNs for a mixed target, composed of a 4-gram and an in-context 2-gram component. This toy setting allows us to capture the entire training process precisely. Specifically, we show that learning progresses through four phases: partial learning of the 4-gram, plateau of induction head learning, emergence of the induction head, and final convergence, showcasing a sharp transition from 4-gram to induction head. Our analysis identifies two key drivers of the transition: 1) time-scale separation due to low- and high-order parameter dependencies in self-attention, and 2) speed differences caused by the relative proportions of the two components in the mixed target.
Finally, we conduct a series of experiments, ranging from simple toy models to real-world natural language training tasks, to validate our theoretical insights.
The detailed discussion of related works is deferred to Appendix A.
@@ 2 Preliminaries
Notations. For $k\in\mathbb{N}^{+}$, let $[k]=\{1,2,\dots,k\}$. For a vector $v$ and $1\leq p\leq\infty$, we denote by $\left\|v\right\|_{p}$ the $\ell_{p}$ norm of $v$. For a matrix $A=(a_{i,j})$, we denote by $\|A\|$, $\|A\|_{F}$ the spectral and Frobenius norms, respectively; let $\|A\|_{1,1}=\sum_{i,j}|a_{i,j}|$. For an event $S$, we define $\mathbb{I}\{S\}=1$ if $S$ is true, and 00 otherwise. We use ${\rm sm}(\cdot)$ to denote the softmax function. We use standard big-O notations $\mathcal{O},\Omega,\Theta$ to hide absolute positive constants, and use $\tilde{\mathcal{O}},\tilde{\Omega},\tilde{\Theta}$ to further hide logarithmic constants.
Sequence modeling. Given a sequence of tokens $(x_{1},x_{2},x_{3},\dots)$ with each token lying in $\mathbb{R}^{d}$, let $X_{L}=(x_{1},x_{2},\dots,x_{L})\in\mathbb{R}^{d\times L}$ and $X_{m:n}=(x_{m}^{\top},x_{m+1}^{\top},\dots,x_{n}^{\top})^{\top}\in\mathbb{R}^{(n-m+1)d}$. Given $A=(a_{1},\cdots a_{n})\in\mathbb{R}^{m\times n}$, we denote $(a_{s})_{s=i}^{j}=(a_{i},\cdots,a_{j})\in\mathbb{R}^{m\times(j-i+1)}$. Then, we consider the next-token prediction task: predict $x_{L+1}$ using $X_{L}=(x_{1},x_{2},\dots,x_{L})$.
@@
Transformer is designed to more efficiently capture long-range dependencies in sequence modeling (Vaswani et al., 2017). Specifically, given an $L$-token input sequence $X=(x_{1},\cdots,x_{L})\in\mathbb{R}^{d\times L}$, an $U$-layer transformer TF processes it as follows. First, each input token is embedded into a higher-dimensional space through an embedding layer with $W_{E}\in\mathbb{R}^{D\times d},b_{E}\in\mathbb{R}^{D}$:
$x_{s}^{(0)}=W_{E}x_{s}+b_{E},\ s\in[L].$
Next, the $U$-layer SA-FFNblocks process the embedded sequence $X^{(0)}=(x_{1}^{(0)},\cdots,x_{L}^{(0)})$ as follows, and the output of the final layer is taken as the output sequence ${\textnormal{{TF}}}(X)=X^{(U)}\in\mathbb{R}^{D\times L}$:
@@
Here, $\textnormal{{FFN}}^{(u)}$ denotes a (token-wise) two-layer FFN of width $M$, and ${\textnormal{{SA}}}^{(u)}$ represents the $H$-head self-attention operation. Specifically, when applied to a sequence $Z=(z_{1},\cdots,z_{L})\in\mathbb{R}^{D\times L}$, ${\textnormal{{SA}}}^{(u)}$ operates it as follows:
$\begin{gathered}\!\!{\textnormal{{SA}}}^{(u)}(Z)=W_{O}^{(u)}\sum_{h=1}^{H}{\textnormal{{SA}}}^{(u,h)}(Z),\quad{\textnormal{{SA}}}^{(u,h)}(Z)=\\ \!\left(W_{V}^{(u,h)}Z\right)\!{\rm softmax}\!\left(\left\langle\!W_{Q}^{(u,h)}\!Z,W_{K}^{(u,h)}\!Z\!\right\rangle\!+\!R^{(u,h)}\!\right)\!,\end{gathered}$ (2)
@@
Relative positional encoding (RPE). Throughout this paper, we focus on the Alibi RPE (Press et al., 2022), where $R_{ij}^{(u,h)}$ exhibit a Toeplitz structure, i.e., $R_{ij}^{(u,h)}=\phi(i-j;p^{(u,h)})$ for $i,j\in[L]$. Here, $p^{(u,h)}$’s are learnable parameters and $\phi(\cdot;p)$ has the following form:
@@
Note that we adopt the Alibi RPE only for simplicity and our results can be extended to other additive RPEs, such as T5 (Raffel et al., 2020), and KERPLE (Chi et al., 2022). However, extending our analysis to the popular rotary RPE (Su et al., 2024) may be nontrivial, and we leave this for future work.
@@ 3.1 Vanilla Induction Heads
[…]
@@
This model retrieves previous tokens $x_{s-1}$’s that are similar to the current token $x_{L}$ based on a dot-product similarity, and then copies and pastes $x_{s-1}$’s subsequent token $x_{s}$ as the current prediction $x_{L+1}$. For example, in Figure 1, the current token $x_{L}$ is D, and the model retrieves previous tokens similar to D, copying and pasting its subsequent token urs as the prediction.
@@
Let $\textnormal{{IH}}_{2}$ satisfy Eq. (4). Then exists an absolute constant $C>0$ and a two-layer single-head transformer TF (without FFNs), with $D=2d$, $W_{K}^{(1,1)}=W_{Q}^{(1,1)}=0$, $p^{(2,1)}=0$, and $\|W_{K}^{(2,1)}\|,\|W_{Q}^{(2,1)}\|\leq\mathcal{O}(1,\|W^{\star}\|_{F})$, such that
@@
This theorem shows that single head suffices to approximate the vanilla induction head and moreover, the approximation efficiency is independent of the sequence length. The proof is provided in Appendix C.1, offering the following insights into how two-layer single-head transformers without FFNs implement vanilla induction heads:
  • The first layer aggregates local tokens and outputs $z_{s}=[x_{s-1},x_{s}]$ for the $s$-th token ($2\leq s\leq L$). This is achieved by using SA with only RPE (no DP). Specifically, RPE allows SA to capture the preceding token via $x_{s-1}=\sum_{j\geq 1}x_{s-j}\rho(j)$ for each token $x_{s}$, where $\rho(\cdot)=\mathbb{I}\{\cdot=1\}$. Hence, DP in this layer is not essential and can be omitted.
  • The second layer extracts the relevant tokens using DP similarity. First, DP computes the similarity $\left\langle W_{Q}z_{L},W_{K}z_{s}\right\rangle=x_{L}^{\top}W^{\star}x_{s-1}$, where $z_{L}=[x_{L-1},x_{L}]$ and $z_{s}=[x_{s-1},x_{s}]$ represent the hidden tokens output by the first layer. This similarity measure enables SA to identify tokens that match $x_{L}$. Subsequently, the value $W_{V}z_{s}$ extracts $x_{s}$, effectively copying the subsequent token of $x_{s-1}$ and using it as the current prediction. In this layer, RPE is not necessary and can be omitted.
@@
Our theoretical analysis is consistent with the experimental observations reported in Elhage et al. (2021). Specifically, the experiments there demonstrate that SA in the first layer attends to adjacent tokens, while SA in the second layer retrieves information related to the current token. Our analysis identifies components responsible for these two operations, and reveals that single-head transformers suffice to perform them efficiently. Furthermore, we validate our theoretical construction through a probing experiment, presented in Figure 3 in Appendix B.1.
@@ 3.2 Generalized Induction Heads: In-context $n$-gram and Generic Similarity
$\begin{gathered}\textnormal{{IH}}_{n}(X_{L})=\sum_{s=n}^{L-1}x_{s}\pi_{s},\\ \pi_{s}={\rm sm}\Big{(}\big{(}X_{L-n+2:L}^{\top}W^{\star}X_{\nu-n+1:\nu-1}\big{)}_{\nu=n}^{L-1}\Big{)}_{\nu=s},\end{gathered}$ (6)
where the patch $X_{\nu-n+1:\nu-1}$ and $X_{L-n+2:L}$ incorporate local information near the previous token $x_{\nu-1}$ and the current $x_{L}$, respectively. This induction head operates based on the similarity between pairs: $(X_{s-n+1:s-1};X_{L-n+2:L})$ for $s=n,\dots,L-1$. Notably, $n$ is typically small when extracting local semantics, so we assume $n\leq 100$ in Eq. (6). [^1] For a given context $X_{1:L}$, Eq. (6) predicts by considering all past $n$-gram pairs $x_{\nu}|X_{\nu-n+1:\nu-1}$ (for $\nu\leq L-1$), making it a context-dependent statistic, and is therefore called in-context $n$-gram. Additionally, note that the $n$ used in classic $n$-grams is typical small (e.g., $2,3,4$), so we make a reasonable assumption that $n\leq 100$.
Integrating richer local information facilitates more accurate information retrieval. The model (6) retrieves previous $(n-1)$-token patch that are similar to the current $(n-1)$-token patch, thereby generalizing the vanilla induction head (4), which considers only single-token retrieval ($n=2$ in Eq. (6)). For example, as depicted in Figure 1, if the current local information is The D (comprising two tokens), and prior local information such as Mr D and Mrs D is identified as similar to The D, transformer would copy and paste their subsequent token, urs, as the prediction.
@@
Let $\textnormal{{IH}}_{n}$ satisfy Eq. (6). Then, for any $H\in\mathbb{N}^{+}$ and rate $q\in\mathbb{N}^{+}$, there exists a constant $C_{n,q}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ (without FFNs), with $D=nd$, such that:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq\left(\frac{C_{n,q}}{H}\right)^{q},$
where $C_{n,q}=\mathcal{O}(nq^{2})$.
This theorem demonstrates that two-layer multi-head transformers, even without FFNs, can efficiently implement the generalized induction head (6). Notably, the approximation error scales as $\mathcal{O}(H^{-q})$, where $q$ can be arbitrarily large. Moreover, for a fixed $q$, $H\geq\Omega(n)$ is sufficient to ensure a good approximation. The proof of this theorem is provided in Appendix C.2. Additionally, the necessity of the required number of heads $H$ and embedding dimension $D$ is experimentally verified in Figure 4 in Appendix B.
@@
$\begin{gathered}\textnormal{{GIH}}_{n}(X_{L})=\sum_{s=n}^{L-1}x_{s}\pi_{s},\\ \ \pi_{s}={\rm sm}\Big{(}\big{(}g\big{(}X_{L-n+2:L};X_{\nu-n+1:\nu-1}\big{)}\big{)}_{\nu=n}^{L-1}\Big{)}_{\nu=s},\end{gathered}$ (7)
@@
Let $\textnormal{{GIH}}_{n}$ satisfy Eq. (7). Suppose the similarity function $g$ is $\alpha$-well-behaved (see Definition C.7). Then there exist two absolute constants $A,B>0$ (only depending on the properties of $g$) such that: for any $H,M\in\mathbb{N}^{+}$ and rate $q\in\mathbb{N}^{+}$, there exists a constant $C_{n,q}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ with FFNs of width $M$, such that
$\left|\!\left|\!\left|\textnormal{{GIH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,2}\leq A\left(\frac{C_{n,q}}{H}\right)^{q}+B\frac{L^{1/(1+2\alpha)}}{M^{\alpha/(1+3\alpha)}},$
where $C_{n,q}=\mathcal{O}(nq^{2})$.
@@
The role of FFNs. In contrast to Theorem 3.3, transformer models in Theorem 3.4 include FFNs. These FFN layers are used to approximate the similarity function $g$. Specifically, we consider the proper orthogonal decomposition (POD) of $g$, which can be viewed as an extension of the matrix singular value decomposition (SVD) applied to functions of two variables. For $g:\mathcal{I}\times\mathcal{I}\to\mathbb{R}$, its POD is $g(u,v)=\sum_{k=1}^{\infty}\sigma_{k}\phi_{k}(u)\psi_{k}(v)$, where $\phi_{k},\psi_{k}$ are orthonormal bases for $L^{2}(\mathcal{I})$ (see Appendix E for details). Intuitively, the FFN in the first layer is used to efficiently approximate $K$ bases ($\phi_{i}$’s and $\psi_{i}$’s). Then, in the second layer, DP in SA can approximately reconstruct $g$ by using the truncated sum $g(u,v)\approx\sum_{k=1}^{K}\sigma_{k}\phi_{k}(u)\psi_{k}(v)$. The complete proof is deferred to Appendix C.3.